May 31, 2011 Readout

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET 8542553 and EVO session
  • JLab Phone: in CC F326 is 757-269-6337 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947
  • Phone:
    • +1-866-740-1260 : US and Canada
    • +1-303-248-0285 : International
    • then use participant code: 3421244# (the # is needed when using the phone)
    • OR from type access code 3421244 into "join a meeting" without the # (you need java plugin)

Items for followup from previous meeting

  1. Check if connections for cooling fluid can be bent radially outward to minimize affect on photons in the gap. Yes. Simulations of impact needed.
  2. Need to get a firm number for the power dissipation in the monitoring control board (in case it is placed inside the housing box) Less than 250 microW/ board.
  3. SiPM: check cross talk between cells (Yi and Carl)

Documents to Review

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. USM Update
  4. SiPM update
  5. Electrical/cooling (Elton)
  6. Geometry and light guides
  7. Simulations


Attending:Tim, David, Beni, Eugene, Elton, Yi (JLab); Christina, George, Pavlos (Athens); Alam, Serguei (USM); Andrei (UofR)

  1. Announcements
  2. Review of Action Items
  3. Monitoring
    • Several exchanges with Kappos last week. Considering the use of flex-boards instead of ribbon cable. Factory in England, but should be practical. The flexible connections would be used for bus and hard PCBs for LED boards as before.
    • New LEDs received. New measurements in progress. Report at next meeting
    • Need to determine distance between boards to order cables. Elton: should wait till connection between LED and light guide in case this has an effect on geometry.
  4. USM Update
    1. Populated board containing pre-amps and multiplexors.
      • Measured signals in response to static inputs, to optimize parameters of pre-amp including output impedance and gain.
      • Eugene: What is the distance between the sensor and the pre-amp. After some discussion, Serguei clarified the distance to be about 10 cm. The scope traces in the attached file were measured with 4-5 cm long wire leads.
    2. Dark Box
      • Adding AC system to stabilize both temperature and humidity in test room.
  5. SiPM update
    • SiPM and electronics was tested with no degradation up to 0.5 T, as expected.
    • Twenty SiPM arrays have been reserved for shipment to UofR (ten units) and USM (ten units). Will prepare for shipment.
  6. Electrical/cooling (Elton)
    1. Fernando is testing an electronic unit (3 pre-amps and one summing circuit). The board is working well, except for a 24 ns rise time that needs to be investigated.
    2. Jack will use the same board, which also contains the bias/temp stabilization circuit, to test the feedback loop.
    3. Mechanical parts for cooling tests were ordered and delivery expected end of May. Need to check on status.
  7. Geometry and light guides
    1. Drawings are being made of the new phi-symmetric light guide configuration.
    2. Prototype light guides from Chile are not yet ready, having some difficulties with cutting the two machined light guides apart. They will also need to be polished once apart.
  8. Simulations
    1. Regular meetings on Thu at 11:00 for a small group to specifically address the effect of segmentation on reconstruction.